a manufacturer purchases two machines A and B .the probability that A will last five years is 4/5 and the probability that B will last five years is 3/4 find the probability that atleast one machine will last for five years
Answer by stanbon(26259)

a manufacturer purchases two machines A and B .the probability that A will last five years is 4/5 and the probability that B will last five years is 3/4 find the probability that atleast one machine will last for five years
------------------
Find the probability that neither will last 5 years:
= (1/5)(1/4) = 1/20
------------------------------
Your Problem:
P(at least will last) = 1 - P(neither will last)
= 1 = 1/20 = 19/20
=============================
Cheers,
Stan H.